# Do Robot Snakes Dream like Electric Sheep? Investigating the Effects of Architectural Inductive Biases on Hallucination

### v2

@@ Abstract
The growth in prominence of large language models (LLMs) in everyday life can be largely attributed to their generative abilities, but some of this is also due to the risks and costs associated with their use. On one front is their tendency to hallucinate false or misleading information, limiting their reliability. On another is the increasing focus on the computational limitations associated with traditional self-attention based LLMs, which has brought about new alternatives, in particular recurrent models, meant to overcome them. However, it remains rare to consider these two concerns simultaneously. Do changes in architecture exacerbate/alleviate existing concerns about hallucinations? Do they affect how and where they occur? Through an extensive evaluation, we study how these architecture-based inductive biases affect the propensity to hallucinate. Although hallucination remains a general phenomenon not limited to specific architectures, the situations in which they occur and the ease with which specific types of hallucinations can be induced can differ significantly depending on the model architecture. These findings highlight the need for better understanding both these problems in conjunction with each other, and consider how to design more universal techniques for handling hallucinations.
@@
[…]
@@ 1 Introduction
Large language models (LLMs) have rapidly emerged as a every-day tool in modern life (OpenAI, 2024), with many relying on their abilities to accomplish a variety of specific tasks. However, this opened up concerns relating to their propensity to hallucinate (Huang et al., 2023), with no concrete reasons for this behaviour (Dziri et al., 2022b; Rawte et al., 2023; Chen et al., 2024), hindering the ability to directly train LLMs that are consistently factual or able to explain themselves through their factual knowledge (Madsen et al., 2024a).
@@
    Evaluating the effects of scale and instruction-tuning and how these changes affect performance, we observe that though factuality hallucinations remain consistently dependent on model size, recurrent/hybrid architectures are often more faithful at smaller sizes and observe significantly fewer faithfulness benefits from instruction-tuning and scaling compared to self-attention models.
@@ 2 Related Work
Though various specific definitions exist, hallucination can broadly refer to "the generation of nonsensical information unfaithful to the provided source content" (Ji et al., 2023a). While this topic has become particularly important in order to ensure the safe and responsible use of language models, both classifying and quantifying the hallucinations is challenging. In particular, it is difficult to ascertain if the divergence occurs because of specific data heuristics (Lebret et al., 2016; Wiseman et al., 2017) or because of the innate lack of similarity between pre-training and downstream tasks (Rashkin et al., 2021), while measuring for hallucinations automatically introduces various biases (Reiter, 2018; Tian et al., 2020; Ganguli et al., 2022) that may not fully capture the scope of the errors. However, a variety of task-specific benchmarks (Li et al., 2020; Pagnoni et al., 2021; Zhou et al., 2021; Santhanam et al., 2022) have shown various LLMs to struggle with factual inconsistencies, highlighting a need to render them safer for every-day use.
@@
General large language models architecture can be broadly thought to be composed of two components: token mixers which serves to model transformations between time steps (such as attention and recurrent layers) and channel mixers, such as multi-layer perceptrons and mixtures-of-experts (MoE) (Jacobs et al., 1991; Shazeer et al., 2017), which allow communication between different channels within a single time step. Accordingly, token mixing often forms a computational bottleneck in terms of time complexity while channel mixers consist of a memory bottleneck. Though the contemporary standard for token mixing remains self-attention, alternatives are becoming increasingly common as they begin to display more promise. These include the use of linear attention (Katharopoulos et al., 2020) to compensate for the quadratic memory complexity of vanilla self-attention, to new recurrent models (Gu and Dao, 2024; Team, 2024b) that function as a linear recurrent neural network but can process all elements of a sequence in parallel, and well as hybrid mixtures of recurrent mechanisms and attention (AI21, 2024; Dao and Gu, 2024) that have emerged as a meaningful competitor.
@@
Despite the growing research in new architectural components, whether or not how they affect hallucination has yet to be studied. While some works (Madsen et al., 2024b; Hu et al., 2024; Schimanski et al., 2024) have proposed modifications to either the learning/generation pipeline as a way of reducing hallucinations, explicit proposals for hallucination reduction through structural modification have yet to be suggested. Additionally, though some works (Elhage et al., 2021; Fu et al., 2023; Lutati et al., 2023; Poli et al., 2024) have demonstrated self-attention to aptly solve synthetic tasks that form an essential component of language modelling, these fail to remain faithful on more realistic datasets, punctuating a major limitation of existing LLMs. Finally, though tangential work demonstrates that recurrent models may suffer from issues with information retention (Vardasbi et al., 2023), formally defining a link with hallucination remains necessary. These issues lead to a lack of clarity on this front, accentuating the need for a formal investigation comparing hallucination alongside architectural paradigms. In particular, the choice of token mixer is important, as this is the component that governs how information is shared between different elements of the sequence to form a holistic representation. As such, an effective token mixer will adequately enable enough information from the context to propagate forward during the generation phase (Olsson et al., 2022; Arora et al., 2024), potentially enabling models to hallucinate less.
@@ 3 Background
Multi-head attention, a variant of self-attention popularized by the Transformer (Vaswani et al., 2017), has become the dominant variant used in LLMs such as LLaMA (Touvron et al., 2023) and Gemma (Team, 2024a). Additional variants, such as multi-query attention from Falcon (Almazrouei et al., 2023) and grouped-query attention from Mistral (Jiang et al., 2023), adapt multi-head attention but remain build upon the same underlying principle of self-attention. Consequently, such methods are considered to fall under the family of self-attention token mixers.
From Transformers to Recurrent LLMs.
One of the original concerns of Transformers was the quadratic complexity of attention, leading to large amounts of work focused on improving this bound as it could directly affect the ability to learn from long sequences. Instead of directly modifying attention (Katharopoulos et al., 2020; Kitaev et al., 2020; Choromanski et al., 2021), Gu et al. (2020) motivated a novel paradigm using state-space models (SSMs) from control theory. SSMs map an input $x(t)\in\mathbb{R}^{d}$ to an intermediate state $h(t)\in\mathbb{R}^{n}$ that is then projected to an output $y(t)\in\mathbb{R}^{d}$:
@@
where $\bm{A}$, $\bm{B}$, $\bm{C}$ and $\bm{D}$ are all trainable parameters. Gu et al. (2021) use this paradigm to define a recurrent model to work on discrete signals, in which case the input can be regarded as discretized data sampled from a continuous signal with a step size $\Delta$, for which the corresponding SSM is defined by:
@@
where $\overline{\bm{K}}$ is the SSM kernel. As $\bm{y}$ can be computed in $O(L\log L)$ with a Fast Fourier Transform (Cormen et al., 2009), the entire output can be computed in tandem based on the input, given the matrices that parameterize the system. Furthermore, setting $\bm{A}$ as a Hurwitz matrix, SSMs can preserve long-term dependency information and enable them to capture long-dependence information, overcoming a long-standing issue (Bengio et al., 1994) with prior recurrent models (Hochreiter and Schmidhuber, 1997; Cho et al., 2014).
Future works then modify this general structure for additional properties; Mamba (Gu and Dao, 2024) use input-dependent $\bm{B}$ and $\bm{C}$, enabling for the model to filter information. LRU/Hawk (Orvieto et al., 2023; De et al., 2024) remove the discretization step from a continuous signal and instead learn discrete matrices for $\bm{A}$, $\bm{B}$, $\bm{C}$ and $\bm{D}$ directly. RWKV models (Peng et al., 2023, 2024) use a novel WKV token mixing operator that acts like an RNN. However, the token mixing paradigm for all such models shares one commonality, namely the use of hidden states from which token-level information must pass through in order to reach another token further in the future. Further interest in this topic has led to the creation of hybrid mixtures of recurrence and attention (Dao and Gu, 2024; De et al., 2024; AI21, 2024) to form new classes of models meant to capture the benefits of both.
Different time-mixing modules attempt to encode the history of tokens in a fixed representation. Recurrent architectures maintain this fixed representation and learn to update it during pre-training. Despite the variety of mechanisms and improvements to the representation according to the design of the module, they can fall short of attention (Vardasbi et al., 2023; Jelassi et al., 2024). One reason is that the amount of information that can be encoded (irrespective of length) is constrained by the size of hidden state (Arora et al., 2024). However, available architectures all converge during training on their respective pre-training and achieve comparable metrics. This then suggests that each architecture learns different aspects of the sequences that align with its own inductive biases. Self-attention architectures have the luxury to look at long range dependencies (albeit being costly). On the other hand, recurrent structures are constrained to make the predictions based on the recent context (which in practice can be long, but remains shorter than what self-attention can attend to). But the expressivity of the hidden representation as modeled by the time-mixers directly affects its performance on tasks (Sun et al., 2024) and the prior knowledge in the parameters of the architectures can be predisposed to favor the prediction of certain tokens. We model this effect on tasks under the lens of hallucination: faithfulness and factuality. Faithfulness expects the model to follow an instruction (that is largely recent), while factuality assumes the use of token-to-token relations learned during pre-training for a broader contextual representation. Following this, we ask two questions: (1) Do self-attention and recurrent architectures hallucinate differently? (2) Does instruction-tuning and/or size benefit the different classes of LLMs equally?
@@ 4 Experiments and Results
For our comparison, we use various models (Table 2) that vary in size and architecture, with a particular focus on different time-mixing methods. These include self-attention models such as Pythia, LLaMA2/3 (Touvron et al., 2023), Falcon-7B (Almazrouei et al., 2023), Mistral-7B (Jiang et al., 2023), Gemma1/2 (Team, 2024a) and Mixtral-8x7B (Jiang et al., 2024). We also test recurrent models including Mamba (Gu and Dao, 2024; Dao and Gu, 2024), FalconMamba-7B and RWKV (Finch) (Peng et al., 2023), as well as hybrid models like RecurrentGemma (Team, 2024b) and Jamba (AI21, 2024). For models with instruction-tuned versions, we use both the base and instruction-tuned variant. Additional details are available in Appendix A.

### v5

@@ Abstract
The growth in prominence of large language models (LLMs) in everyday life can be largely attributed to their generative abilities, yet some of this is also owed to the risks and costs associated with their use. On one front is their tendency to hallucinate false or misleading information, limiting their reliability. On another is the increasing focus on the computational limitations associated with traditional self-attention based LLMs, which has brought about new alternatives, in particular recurrent models, meant to overcome them. Yet it remains uncommon to consider these two concerns simultaneously. Do changes in architecture exacerbate/alleviate existing concerns about hallucinations? Do they affect how and where they occur? Through an extensive evaluation, we study how these architecture-based inductive biases affect the propensity to hallucinate. While hallucination remains a general phenomenon not limited to specific architectures, the situations in which they occur and the ease with which specific types of hallucinations can be induced can significantly differ based on the model architecture. These findings highlight the need for better understanding both these problems in conjunction with each other, as well as consider how to design more universal techniques for handling hallucinations.
@@
Jerry Huang Chandar Research Lab Mila & Université de Montréal                        Prasanna Parthasarathithanks: Corresponding author: pp1403@gmail.com Noah’s Ark Lab                        Mehdi Rezagholizadehthanks: Work done while at Noah’s Ark Lab. Advanced Micro Devices
Boxing Chen Noah’s Ark Lab                        Sarath Chandar Chandar Research Lab Mila & Polytechnique Montréal CIFAR AI Chair
@@ 1 Introduction
Large language models (LLMs) have rapidly emerged as a every-day tool in modern life (OpenAI, 2024), with many relying on their abilities to accomplish a variety of specific tasks. However, this opened up concerns relating to their propensity to hallucinate (Huang et al., 2023), with no concrete reasons for this behaviour (Dziri et al., 2022b; Rawte et al., 2023; Chen et al., 2024b), hindering the ability to directly train LLMs that are consistently factual or able to explain themselves through their factual knowledge (Madsen et al., 2024a; Prato et al., 2023, 2024; Huang et al., 2024).
@@
    Evaluating the effects of scale and instruction-tuning, we observe that though factuality remains dependent on model size, recurrent/hybrid architectures are often more faithful at smaller sizes and observe significantly fewer faithfulness benefits from instruction-tuning and scaling compared to self-attention models.
@@ 2 Related Work
Hallucination broadly refers to when LLMs generate information that does not directly follow from the context, such as nonsensical or irrelevant answers to questions (Ji et al., 2023a). While it has grown in importance due to its direct relationship with ensuring the safe and responsible use of LLMs, both classifying and quantifying the hallucinations is challenging. In particular, it is difficult to ascertain if the divergence occurs because of specific data heuristics (Lebret et al., 2016; Wiseman et al., 2017) or because of the innate lack of similarity between pre-training and downstream tasks (Rashkin et al., 2021), while measuring for hallucinations automatically introduces various biases (Reiter, 2018; Tian et al., 2020; Ganguli et al., 2022) that may not fully capture the scope of the errors. However, a variety of task-specific benchmarks (Li et al., 2020; Pagnoni et al., 2021; Zhou et al., 2021; Santhanam et al., 2022) have shown various LLMs to struggle with factual inconsistencies, highlighting a need to render them safer for every-day use.
@@
General LLM architecture can be broadly thought to be composed of two components: token mixers which serves to model transformations between time steps (such as attention and recurrent layers) and channel mixers, such as multi-layer perceptrons and mixtures-of-experts (MoE) (Jacobs et al., 1991; Shazeer et al., 2017), which allow communication between different channels within a single time step. Accordingly, token mixing often forms a computational bottleneck in terms of time complexity while channel mixers consist of a memory bottleneck. Though the contemporary standard for token mixing remains self-attention, alternatives are becoming increasingly common as they begin to display more promise. These include the use of linear attention (Katharopoulos et al., 2020) to compensate for the quadratic memory complexity of vanilla self-attention, to new recurrent models (Gu and Dao, 2024; Gemma Team, 2024b) that function as a linear recurrent neural network but can process all elements of a sequence in parallel, and well as hybrid mixtures of recurrent mechanisms and attention (AI21, 2024; Dao and Gu, 2024) that have emerged as a meaningful competitor.
@@
Despite growing research in new architectural components, their effects on hallucination have yet to be studied. While some works (Madsen et al., 2024b; Hu et al., 2024; Schimanski et al., 2024) have proposed modifications to either the learning/generation pipeline as a way of reducing hallucinations, proposals for hallucination reduction through structural modification have yet to be suggested. Additionally, though some works (Elhage et al., 2021; Fu et al., 2023; Lutati et al., 2023; Poli et al., 2024) have demonstrated self-attention to aptly solve synthetic tasks that form an essential component of language modeling, these fail to remain faithful on more realistic datasets, punctuating a major limitation of existing LLMs. Finally, though tangential work demonstrates that recurrent models may suffer from issues with information retention (Vardasbi et al., 2023), formally defining a link with hallucination remains necessary. These issues lead to a lack of clarity on this front, accentuating the need for a formal investigation comparing hallucination alongside architectural paradigms. In particular, the choice of token mixer is important, as this is the component that governs how information is shared between different elements of the sequence to form a holistic representation. As such, an effective token mixer will adequately enable enough information from the context to propagate forward during the generation phase (Olsson et al., 2022; Arora et al., 2024), potentially enabling models to hallucinate less.
@@ 3 Background
Multi-head attention, a variant of self-attention popularized by the Transformer (Vaswani et al., 2017), has become the dominant variant used in LLMs such as LLaMA (LLaMA Team, 2024) and Gemma (Gemma Team, 2024a). Additional variants, such as multi-query attention from Falcon (Almazrouei et al., 2023) and grouped-query attention from Mistral (Jiang et al., 2023), adapt multi-head attention but remain build upon the same underlying principle of self-attention. Consequently, such methods are considered to fall under the family of self-attention token mixers.
Recurrent LLMs.
A concern for Transformers is the quadratic complexity of attention with sequence length, leading a focused on improving this bound as it directly affects the ability to learn from long sequences. Instead of directly modifying attention (Katharopoulos et al., 2020; Kitaev et al., 2020; Choromanski et al., 2021; Zeng et al., 2025), Gu et al. (2020) motivated a novel paradigm using state-space models (SSMs) from control theory. SSMs map an input $x(t)\in\mathbb{R}^{d}$ to an intermediate state $h(t)\in\mathbb{R}^{n}$ that is then projected to an output $y(t)\in\mathbb{R}^{d}$:
@@
where $\bm{A}$, $\bm{B}$, $\bm{C}$ and $\bm{D}$ are trainable parameters and $h^{\prime}(t)$ represents the rate at which $h(t)$ changes. Gu et al. (2021) use this paradigm to define a recurrent model to work on discrete signals, in which case the input can be regarded as discretized data sampled from a continuous signal with a step size $\Delta$, for which the corresponding SSM is defined by:
@@
where $\overline{\bm{K}}$ is the SSM kernel. As $\bm{y}$ can be computed in $O(L\log L)$ with a Fast Fourier Transform (Cormen et al., 2009), the entire output can be computed in tandem, given the matrices that parameterize the system. Furthermore, setting $\bm{A}$ as a Hurwitz matrix, SSMs can preserve long-term information, overcoming a long-standing issue (Bengio et al., 1994) with prior recurrent models (Hochreiter and Schmidhuber, 1997; Cho et al., 2014).
Further works have modified this structure; Gu and Dao (2024) use input-dependent $\bm{B}$ and $\bm{C}$ information filtering. LRU/Hawk (Orvieto et al., 2023; De et al., 2024) remove the discretization step from a continuous signal and instead learn discrete matrices for $\bm{A}$, $\bm{B}$, $\bm{C}$ and $\bm{D}$ directly. RWKV (Peng et al., 2023, 2024) uses a novel WKV sequence mixing operator that acts like an RNN. However, these models all share the use of hidden states from which token-level information must pass through in order to interact with future tokens. This has also led to the creation of hybrid mixtures of recurrence and attention (Dao and Gu, 2024; De et al., 2024; AI21, 2024) to form new classes of models meant to capture the benefits of both.
Different sequence (or time) mixing modules attempt to encode the history of tokens in a fixed representation. Recurrent architectures maintain this representation and learn to update it during pre-training, but this can fall short of attention (Vardasbi et al., 2023; Jelassi et al., 2024; Huang, 2025; Wang et al., 2025) as they are constrained by its size (Arora et al., 2024). Meanwhile, attention has a theoretically infinite context window (albeit being costly). As recurrent structures are constrained to make the predictions based on the recent context (which in practice can be long, but remains shorter than what self-attention can attend to), the expressivity of the hidden representation directly affects task performance (Sun et al., 2024) and parametric knowledge can be predisposed to favor the prediction of certain tokens. We model this effect on tasks under the lens of hallucination: faithfulness and factuality. Faithfulness expects the model to follow an instruction (that is largely recent), while factuality assumes the use of token-to-token relations learned during pre-training for a broader contextual representation. Following this, we ask two questions: (1) Do self-attention and recurrent architectures hallucinate differently? (2) Does instruction-tuning and/or size benefit the different classes of LLMs equally?
@@ 4 Experiments and Results
To investigate, we use models that vary in size and architecture, with a particular focus on different sequence mixing methods. These include self-attention models (Pythia, LLaMA2/3 (LLaMA Team, 2024), Falcon (Almazrouei et al., 2023), Mistral (Jiang et al., 2023), Gemma (Gemma Team, 2024a) and Mixtral (Jiang et al., 2024)), recurrent models (Mamba (Gu and Dao, 2024; Dao and Gu, 2024), FalconMamba and RWKV/Finch (Peng et al., 2023)), as well as hybrid models (RecurrentGemma (Gemma Team, 2024b), Jamba (AI21, 2024)). We use base and instruction-tuned variants when available. Additional details are available in Appendix A.